OpenStudy (anonymous):

I think its 2(t+8) ,help? Write an algebraic expression for each word phrase. 2 times the sum of t and 8 A. 2t(t+8) B. 2(t+8) C. 2(t-8) D. (8-2t)

OpenStudy (anonymous):

yes it is 2(t+8)

OpenStudy (anonymous):

Sum of t and 8 = t + 8 2 x sum of t and 8 = 2(t+8)

OpenStudy (anonymous):

ok thnx just makin sure !

OpenStudy (anonymous):

Yep ur correct :)